# Expressivity of Neural Networks with Random Weights and Learned Biases

### v1

@@ 2.2 Recurrent neural networks
This section studies a discrete-time RNN given by
$r_{t}=-\alpha r_{t-1}+\beta\phi(Wr_{t-1}+Bx_{t-1}+b),\quad y_{t}=Cr_{t},$ (2)
@@
where $t$ and $x_{t}$ are as defined for the RNN, $F:U_{z}\times U_{x}\to\mathbb{R}^{s}$ is continuous, and $Q\in\mathbb{R}^{l\times s}$. Because we build from the classic universal approximation results, we must be working with functions operating on compact sets. To guarantee that this will be the case we must make several more assumptions about the dynamical system. First, $U_{z}\subset\mathbb{R}^{s}$ is assumed to be a compact invariant set of the dynamical system: if the system is in $U_{z}$ it remains there for all $t$ for all inputs in $U_{x}$. Second, we assume that the dynamical system is well-defined on a slightly larger compact set $\tilde{U}_{z}\times U_{x}$, where $\tilde{U}_{z}=\{z_{0}+c_{0}:z_{0}\in U_{z},\>||c_{0}||<c\}$ for some $c>0$, with $\tilde{U}_{z}\supset U_{z}$. Letting $p_{R}$ be defined as in Section 2.1, the main result of this section is:
@@
Consider the RNN in Eq.2 with $\phi$ a $\gamma$-bias learning activation, and input, output, and recurrent weight parameters for each hidden unit sampled from $p_{R}$. We can find a hidden-layer width, a bias vector, and a hidden-state initial condition for the RNN such that, with a probability that is arbitrarily close to $1$, the RNN approximates finite trajectories from the dynamical system defined in Eq.3 to below any positive, non-zero, error.
This proof is weaker than the proof given in the previous section in that it shows point-wise convergence, rather than $L^{1}$, convergence. We conjecture that this can be extended straightforwardly but we leave this result to future work.
Proof Intuition: The complete proof is in the Appendix; here we provide intuition. As in Theorem 1 the proof proceeds in two steps. First, the dynamical system is approximated by an RNN, $\mathcal{R}_{1}$, using universal approximation theory for RNNs (see e.g., Schäfer and Zimmermann (2006)). $\mathcal{R}_{1}$ is then approximated by a much wider, random RNN, $\mathcal{R}_{2}$, with parameters sampled from $p_{R}$. Analogous to Theorem 1, we show that one can find a sub-network of hidden units in $\mathcal{R}_{2}$ that approximates $\mathcal{R}_{1}$ for very large hidden widths of $\mathcal{R}_{2}$.
@@ 3 Numerical Results
3.1 Multi-task Learning with Bias-learning Feed-forward Networks
We first validated the theory by checking whether a single-hidden-layer, bias-learning FFN could perform digit recognition on the MNIST dataset Deng (2012) increasingly well as its hidden layer was widened. Initial weights were sampled from a uniform distribution on $[-0.1,0.1]$ and then frozen. As expected, the network learned the task and validation accuracy increased with layer width. The largest gains in performance occurred before 5000 units (Fig. 1A).
@@ 3.1 Multi-task Learning with Bias-learning Feed-forward Networks
Figure 1: A. Validation accuracy on MNIST vs layer width. B. Validation accuracy on multiple image classification tasks for bias-only (blue) and fully-trained (orange) networks. In panels A and B, training was on 20 epochs and error bars on 5 runs are omitted because the standard errors are of order $10^{-3}$. C. Top: K-mean clustering of Task Variance (TV) reveals task-specific clusters. Bottom: Spearman correlation between TV and bias vectors (mean across neurons in each cluster). In this and all other figures Adam Kingma and Ba (2014) was used for optimization.
@@
Intuitively, bias learning should allow a single random set of weights to be used to learn multiple tasks by simply optimizing task-specific bias vectors. We confirmed this intuition by training a single-hidden-layer FFN with 32000 hidden units on 7 different tasks: MNIST Deng (2012), KMNIST Clanuwat et al. (2018), Fashion MNIST Xiao et al. (2017), Ethiopic-MNIST, Vai-MNIST, and Osmanya-MNIST from Afro-MNIST Wu et al. (2020), and Kannada-MNIST Prabhu (2019). All tasks involved classifying 28$\times$28 grayscale images into 10 classes. The random weights were fixed across tasks while different biases were learned. We compared bias learning against a fully-trained neural network with the same size and architecture (Fig. 1B). We found that the bias-only network achieved similar performance to the fully-trained network on most tasks (only significantly worse on KMNIST). A crucial difference here is that the networks had matched size and architecture, so that the number of trainable parameters in the bias-only network (32 000 parameters) was several orders of magnitude smaller than in the fully-trained case (25 440 010 parameters). Notably, a different set of biases was learned for each task. We conclude that bias-only learning in FFNs is a viable avenue to perform multi-tasking with randomly initialized and fixed weights, but that it requires a much wider hidden layer than fully trained networks.
@@
Next, we investigated the neural mechanism underlying bias learning of multiple tasks. We investigated the task-specific functional organization of the hidden units by estimating single-unit Task Variance (TV) Yang et al. (2019), defined as the variance of a hidden unit activation across the test set for each task. The TV provides a measure of the extent that a given hidden unit contributes to the given task. A unit with high TV in a particular task indicates that its responses vary across stimuli, suggesting that the unit is recruited for solving that task. A unit with high TV in one task and low TV for all other tasks is specialized to one particular task. We clustered the hidden units TV using K-means clustering ($K$ chosen by cross-validation) on the vectors of TVs for each unit and found that distinct functional clusters of hidden units emerged (Fig. 1C). Most units reflected strong task specialization, i.e., they were only used for specific tasks (ex: cluster 3 for KMNIST and cluster 10 for Osmanya). Others were used for many or all tasks (ex: clusters 1 and 8), although a smaller fraction of clusters exhibited such non-selective activation patterns. Overall, we conclude that multi-task bias-learning leads to the emergence of task-specific functional organization.
@@
We finally explored the relationship between the bias of a hidden unit and its task variance (TV). If the neural networks are using biases to shut-off units, analogous to the intuition in our theory work (Section 2.1), then the network units that do not actively participate in a task should be more quiet due to a low bias value learned during training on that particular task. In other words, this intuition would suggest that units should exhibit a correlation between bias and TV, especially in task-specific clusters. In our experiments, all clusters did exhibit the statistical trend of a positive correlation between bias and TV, although to a varying degree across clusters (see numbers at the bottom of Fig. 1C).
3.2 Relationship Between Bias Learning and Masked Learning in FFNs
Figure 2: Comparing bias and masked learning on same weights. A. Learning curves for bias (orange) and mask (black) learning. Inset: we observed a trend where bias-learning achieved roughly $1$ percentage point higher accuracy on MNIST over mask-learning ($0.915\pm 0.0028\mathrm{SD}$ bias vs. $0.905\pm 0.0028\mathrm{SD}$ mask). B. Histograms of hidden unit variances, calculated over $10,000$ MNIST samples, for bias-trained (orange) and mask-trained (black). Histogram counts are log-scaled. C. Scatter plot of hidden unit variances from C but taking only units that are non-zero/not approximately-zero in mask/bias-trained networks; bias-trained on x-axis and mask-trained on y-axis. Plot A is mean±$\pm 1$sd over 4 mask-trained/bias-trained nets. Plot B, and C are both one representative network trained with biases and with masks. Learned parameters were initialized to uniform on $[-0.1,0.1]$, weights to uniform on $[-\frac{1}{\sqrt{d}},\frac{1}{\sqrt{d}}]$; all other hyper-parameters can be found in the codebase.
@@ 3.2 Relationship Between Bias Learning and Masked Learning in FFNs
As our theory shows that bias-learning networks can universally approximate simply by turning units off, we wished to test whether bias learning performs similarly to learning masks, and to what extent solutions learned by these approaches are different from each other. We compared training mask to bias learning on networks with the same random input/output weight matrices. For mask-training, we approximated binary masks with ‘soft’ sigmoid masks by optimizing the sigmoid parameters. The sigmoid was steepened over the course of training to approximately learn a mask, and at test time the slope of the sigmoid was made steep enough to effectively binarize the learned soft mask. We compared masks learned in this fashion with learned biases on single-hidden-layer ReLU networks with $10,000$ hidden units. We observed a trend of bias-training slightly improving upon mask-training (Fig.2A), an expected increase given the added flexibility of biases over masks. Further research is needed to determine if this trend is reliable across datasets and different network parametrizations, and whether there might be scenarios where one style of learning works better or worse.
Next, we compared the solutions found via bias and mask learning. We calculated the variance of each hidden unit across $10,000$ MNIST images, in both the bias and mask-trained paradigms, as a measure of the hidden-layer representation of MNIST. The histogram over hidden units of these variances is plotted for trained masks/biases (black/orange) for the same network weights (Fig.2B). The mask-trained networks found slightly sparser solutions (more 00s, far left histogram bin) and performed the task with lower unit variance values (see middle/right of histogram) compared to the bias-trained networks. We also plotted the hidden unit variances for units in mask and bias-trained networks and computed their correlation (Fig.2C). The scatter plot and correlation were calculated after removing unit variances that were zero in the mask-trained net and sufficiently close to zero (with variance values less than the mean variance divided by 100) in the bias-trained network. Despite the differences observed in the histogram, the correlation coefficient was $0.461\pm 0.022$ (mean $\pm$ SD across $n=4$ networks), suggesting that there is some overlap in the way that these two methods solve MNIST.
3.3 Bias-learning of Dynamical Systems with Recurrent Neural Networks
3.3.1 Autonomous dynamical systems
@@
Figure 3: Learning autonomous dynamical systems. A: Cosine generated by a bias-learning recurrent network with 200 hidden units (dashed orange) and its target (solid black). B. Eigenvalue spectra for the recurrent weight matrix (left) and the Jacobian at the start of training (right, grey squares) and mid-training (right, orange circles), when the network produced a decaying oscillation with period 23.75, close to the target period of 25. Units’ activity approached a fixed point with respect to which the Jacobian was computed. Later in training, the rightmost eigenvalues approached 1 in magnitude and their phases were such that the oscillation had period $\sim 25$. C. Van der Pol oscillator (target in solid black; with dynamics $\ddot{x}=\mu(1-x^{2})\dot{x}-x$, for $\mu=2$) generated by the bias-learning recurrent network with 675 hidden units for a recurrent gain of 1 (dashed orange; see panel D) and a gain of 0.9 (dashed dark orange). Output represents the position of the oscillator, rescaled to [-1, 1]. D. (Left) Sensitivity to distribution of recurrent weights. The fully-trained and bias-learning networks had the same number of learnable parameters (size of hidden layer for fully-trained network was 25). Initial recurrent weight matrix had elements sampled from $(g/\sqrt{m})\mathcal{N}(0,1)$, where $g$ is the gain (Gain recurrent init.). Error bars denote SEM for $n=10$. (Right) Schematics of the fully-trained (top) and bias-learning (bottom) autonomous RNNs. Colored links denote trained weights. In all panels, output matrix elements were $\sim\mathcal{N}(0,1/m^{2})$ and biases $b\sim\mathcal{U}(0,1)$.
@@
We next investigated whether bias learning relied on the statistics of the fixed recurrent weights. In light of Fig. 3B, we thus hypothesized that bias learning would be affected by changes in the weight distribution, because bias learning can only control the derivative. We initialized an i.i.d. Gaussian distributed weight matrix $W_{ij}\sim{\cal N}(0,g^{2}/N)$, where $g$ is referred to as its ‘gain’. We then trained bias-learning networks to generate a van der Pol oscillator (Fig. 3C). We found that bias learning required a large enough gain (at least $g=1$) and failed for $g<1$ (Fig. 3D). This was not purely due to a restricted dynamic range for the network activity since the network was able to reproduce the first peak of the oscillator and then flatlined (Fig. 3C). In contrast, fully-trained networks with the same number of training parameters (Fig. 3D) were not sensitive to the value of the gain at initialization. According to our theory, because ReLU is $\gamma$-parameter bounding, a bias-learning RNN should be able to produce the correct output even at low gain, but this could occur for a prohibitively large hidden layer. This result thus highlights that, when the hidden-layer size is fixed, the initial distribution of weights limits the capability of bias-learning networks.
3.3.2 Non-autonomous dynamical systems
[…]
Figure 4: Learning non-autonomous dynamical systems. A. The outputs of the fully-trained (top) and bias-only (bottom) networks on a window of a Lorenz system unseen during training ($\dot{x}=\sigma(y-x),\dot{y}=x(\rho-z)-y,\dot{z}=xy-\beta z$). The system was generated using Euler’s method with a step size $\Delta t$ of 0.01 from an initialization at (0,1,0), with $\sigma=10$, $\rho=28$, and $\beta=\frac{8}{3}$. Standard deviation error bars were computed over 5 seeds, but are not visible. B. Generalization to sequences longer (4320 time-points) than those trained on. C. Output of the bias-only network diverges from the ground truth signal when using its own outputs as context (starting from the grey line).

### v3

@@ 2.2 Recurrent neural networks
Here, we study a discrete-time RNN given by:
$r_{t}=-\alpha r_{t-1}+\beta\phi(Wr_{t-1}+Bx_{t-1}+b),\quad\hat{y}_{t}=Cr_{t},$ (2)
@@
where $t$ and $x_{t}$ are as defined for the RNN, $F:U_{z}\times U_{x}\to\mathbb{R}^{s}$ is continuous, and $Q\in\mathbb{R}^{l\times s}$. Because we build from the classic universal approximation results, we must be working with functions operating on compact sets. To guarantee that this will be the case we must make several more assumptions about the dynamical system. First, $U_{z}\subset\mathbb{R}^{s}$ is assumed to be a compact invariant set of the dynamical system: if the system is in $U_{z}$ it remains there for all $t$ and for all inputs in $U_{x}$. Second, we assume that the dynamical system is well-defined on a slightly larger compact set $\tilde{U}_{z}\times U_{x}$, where $\tilde{U}_{z}=\{z_{0}+c_{0}:z_{0}\in U_{z},\>||c_{0}||<c\}$ for some $c>0$, with $\tilde{U}_{z}\supset U_{z}$.
To generalize the approximation in $L^{\infty}$ norm in our analysis of FNNs to dynamical systems we consider an infinity norm over finite trajectories: $\sup_{z_{0}\in U_{z},\mathbf{x}_{t}\in U_{\mathbf{x}}}\sum_{t=1}^{T}||y_{t}(z_{0},\mathbf{x}_{t})-\hat{y}_{t}(r_{0},\mathbf{x}_{t})||$, where $\mathbf{x}_{t}\equiv[x_{0},\dots,x_{t-1}]$, $r_{0}\equiv r_{0}(z_{0})$ is a continuous map from $U_{z}$ into the hidden state space of the RNN, and $U_{\mathbf{x}}$ is the $t$-times product space of $U_{x}$. Letting $p_{R}$ be defined as in Section 2.1, the main result of this section is:
@@
Consider the RNN in Eq.2 with $\phi$ a $\gamma$-bias-learning activation, and input, output, and recurrent weight parameters for each hidden unit sampled from $p_{R}$. We can find a hidden-layer width, $m$, a bias vector, and a continuous hidden-state initial condition map $r_{0}:U_{z}\mapsto\mathbb{R}^{m}$ such that, with a probability that is arbitrarily close to $1$, the RNN approximates the dynamical system defined in Eq.3 to below any positive, non-zero, error in the inifinity norm over trajectories.
Proof Intuition: We provide a high-level description here while the detailed proof and a schematic (Fig. E.1) can be found in the Appendix. As in Theorem 1 the proof proceeds in two steps. First, the dynamical system is approximated by an RNN, $\mathcal{R}_{1}$, using universal approximation theory for RNNs (see e.g., Schäfer & Zimmermann (2006)). $\mathcal{R}_{1}$ is then approximated by a much wider, random RNN, $\mathcal{R}_{2}$, with parameters sampled from $p_{R}$. Analogous to Theorem 1, we show that one can find a sub-network of hidden units in $\mathcal{R}_{2}$ that approximates $\mathcal{R}_{1}$ for very large hidden widths of $\mathcal{R}_{2}$..
@@ 3 Numerical results
3.1 Multi-task learning with bias-learned FNNs
We first validated the theory by checking whether a single-hidden-layer bias-learned FNN could perform classification on the Fashion MNIST dataset Deng (2012) increasingly well as its hidden layer was widened. We compared fully-trained networks, matching the number of trained parameters, with bias-learned networks where the frozen weights were sampled from a uniform distribution on $[-0.1,0.1]$ or a zero-mean Gaussian with standard deviation $\frac{1}{\sqrt{d}}$, where $d$ is the input dimension. The networks successfully learned the task and validation error decreased with the number of trained parameters (Fig. 1A) which, for bias-learned networks, is equal to the number of hidden units. The largest gains in performance occurred before $5\times 10^{3}$ units, with bias-learning achieving comparable, but slightly worse, performance compared to fully-trained models. We speculate that this gap in performance for large network sizes is due either to current deep learning training conventions being optimized for weight training, or to bias learning requiring even larger hidden layer sizes to match fully-trained performance. Interestingly, for very low parameter counts bias-learning out-performed fully trained networks (more visible with log-log axis scaling E.2.A). Because weights scale quadratically with layer width, a fully-trained network will have a smaller width than a bias-learned network with the same number of trained parameters (see Fig.E.2.B for the error scaling plot with layer width on the $x$-axis).
@@ 3.1 Multi-task learning with bias-learned FNNs
Figure 1: A. Validation accuracy on fashion MNIST vs. number of trained parameters for fully-trained (blue), bias-learned with uniformly distributed weights (light orange), and bias-learned with Gaussian weights (dark orange) networks. B. Validation accuracy on multiple image classification tasks for bias-learned (orange) and fully-trained (blue) networks. Errors for 5 random seeds are barely visible as the shaded regions in A, and are omitted in B because the standard errors are of order $10^{-3}$. C. Top: K-mean clustering of Task Variance (TV) reveals task-selective clusters (see Fig.E.3 for fully-trained network selectivity). Bottom: Spearman correlation between TV and bias vectors (mean across neurons in each cluster).
@@
Intuitively, bias learning should allow a single random set of weights to be used to learn multiple tasks by simply optimizing task-specific bias vectors. We confirmed this by training a single-hidden-layer FNN with $3.2\times 10^{4}$ hidden units on 7 different tasks: MNIST Deng (2012), KMNIST Clanuwat et al. (2018), Fashion MNIST Xiao et al. (2017), Ethiopic-MNIST, Vai-MNIST, and Osmanya-MNIST from Afro-MNIST Wu et al. (2020), and Kannada-MNIST Prabhu (2019). All tasks involved classifying 28$\times$28 grayscale images into 10 classes. The random weights were fixed across tasks while different biases were learned. We compared bias learning against a fully-trained neural network with the same size and architecture (Fig. 1B). We found that the bias-only network achieved similar performance to the fully-trained network on most tasks (only significantly worse on KMNIST). An important difference here is that the networks had matched size and architecture, so that the number of trainable parameters in the bias-only network ($3.2\times 10^{4}$ parameters) was several orders of magnitude smaller than in the fully-trained case ($\approx 2.5\times 10^{7}$ parameters). Notably, a different set of biases was learned for each task. We conclude that bias-only learning in FNNs could be a viable avenue to perform multi-tasking with randomly initialized and fixed weights, but that it requires a much wider hidden layer than fully trained networks. Lastly, we note that the networks of Fig. 1B were trained with uniformly initialized weights, but that one can achieve similar, or even better performance with different weight initializations (see Fig. E.2C).
@@
Next, we investigated the task-specific responses of hidden units by estimating single-unit Task Variances (TV) Yang et al. (2019), defined as the variance of a hidden unit activation across the test set for each task. The TV provides a measure of the extent that a given hidden unit contributes to the given task: a unit with high TV in one task and low TV for all others is seen as selective for its high-TV task. We clustered the hidden-unit TVs using K-means clustering ($K$ chosen by cross-validation) on the vectors of TVs for each unit and found that distinct clusters of units emerged (Fig. 1C). Some units reflected strong task selectivity (ex: cluster 3 for KMNIST and cluster 10 for Osmanya). Others responded to many, or all, tasks (ex: clusters 1 and 8), although a smaller fraction of clusters exhibited such non-selective activation patterns. Overall, we conclude that multi-task bias learning leads to the emergence of task-specific organization. We note, however, that task selectivity does not necessarily mean task utility: for example, a neuron could have a high variance for a single task but that variance could be picking up on noise and thus not functionally useful. We leave a deeper investigation of the functional significance of task-selectivity to future work.
@@
Finally, we explored the relationship between the bias of a hidden unit and its TV. If the neural networks are using biases to shut-off units, analogous to the intuition in our theory (Section 2.1), then the units that do not actively participate in a task should be quiet due to a low bias value learned during training on that particular task. In other words, this intuition would suggest that units should exhibit a correlation between bias and TV, especially in task-specific clusters. In our experiments, all clusters did exhibit the statistical trend of a positive correlation between bias and TV, although to a varying degree across clusters (see numbers at the bottom of Fig. 1C).
3.2 Relationship between bias learning and mask learning in FNNs
Figure 2: Comparing bias and mask learning on same weights. A. Learning curves for bias (orange) and mask (black) learning on MNIST. Inset: bias learning achieved roughly $1\%$ higher test accuracy over mask learning ($0.934\pm 0.001\mathrm{SD}$ bias vs. $0.919\pm 0.002\mathrm{SD}$ mask). B. Probability ($y$-axis) of the same unit being ON in both the bias-learning and mask-learning networks (orange line). A unit is ‘ON’ in mask learning if it is not masked out, and in bias learning if it has task variance above a given threshold ($x$-axis). Also shown is the probability of a unit being ON in two different training runs for mask-learning (black dashed line), and a null model giving the expected overlap if the probability of a unit being ON in the bias-trained network is independent of whether it is ON in the mask network (see Appendix §C.2 for more details) C. Histograms of hidden unit variances, calculated over $10^{4}$ test set MNIST samples, for bias-trained (orange) and mask-trained (black). Unit variances below $0.1$ are not shown. All curves, and histograms, are means, with shaded regions being $1$SD over $5$ training runs.
@@ 3.2 Relationship between bias learning and mask learning in FNNs
As our theory shows that bias learning networks can universally approximate simply by turning units off, we wished to test whether bias learning performs similarly to learning masks, and to what extent solutions learned by these approaches are different from each other. We compared training mask to bias learning on networks with the same random input/output weight matrices. For mask-training, we approximated binary masks using ‘soft’ sigmoid masks with learned gain parameters (see Methods). The approximation of a discontinuity with a differentiable function was done to allow optimization with gradient descent, a strategy with a history of use in ML Jang et al. (2016) and computational neuroscience Zenke & Ganguli (2018). The sigmoid was steepened over the course of training to approximate the binary mask that was used at test time. We compared masks learned in this fashion with learned biases on single-hidden-layer ReLU networks with $10^{4}$ units. We observed a trend of bias-training slightly improving upon mask-training (Fig.2A), which was expected given that biases can be tuned over a continuous range of values, including 00 and $1$, while masks can only take 00 or $1$. Further research is needed to determine if this trend is reliable across datasets and different network parametrizations, and whether there might be scenarios where one style of learning works better or worse.
Next, we compared the solutions found via bias and mask learning on the same set of randomly initialized weights. We calculated the variance of each hidden unit across $10^{4}$ MNIST test images, in both the bias and mask-trained paradigms, as a measure of hidden layer representation. To investigate whether the same units were active during the task regardless of training style, we looked at which units were ‘ON’ in mask versus bias-trained networks. For mask learning a unit was considered ON if its mask was $1$; for bias learning a unit was ON if its task variance was above a given threshold. We observed that neurons with higher task variances contributed more to solving the task (Fig. E.4A). For a range of thresholds, we calculated the probability that a given unit was ON for both mask and bias learning (Fig.2B). We found that, for low thresholds, this match probability was intermediate between chance (grey line) and the high probability of a match between two mask-learned training runs on the same set of weights (black dashed line). We further noted that, on average, mask learning used $4527\pm 44$ ‘ON’ units (in this section all reported values are mean±$\pm$±SD over $5$ samples) to solve the task. For bias learning, when the task variance threshold was moved from 00 to $0.1$ test accuracy dropped about $1\%$ and the number of ON units went from $10^{4}$ (all units) to $6226\pm 47$. Thus mask learning, found a sparser solution than bias learning. To further investigate the differences and similarities between unmasked units, we plotted the histograms of unit variances above the $0.1$ threshold, and observed that bias learning used higher variances to accomplish the task (Fig.2C). Finally, we found a moderate correlation between the task variances of units that were ON for both mask and bias learning ($0.5030\pm 0.0087$) (Fig E.4C). In summary, we observe that, relative to mask learning, bias learning finds a different, but overlapping, solution to MNIST.
3.3 Bias learning autonomous dynamical systems with RNNs
@@
Figure 3: Learning autonomous dynamical systems. A. Cosine generated by a bias-learning RNN (dashed orange) and its target (solid black). B. Eigenvalue spectra for the recurrent weights (left) and the Jacobian at the start of training (right, grey squares) and mid-training (right, orange circles), when the network produced a decaying oscillation with period 23.75, close to the target period of 25. Neural activity then approached a fixed point with respect to which the Jacobian was computed. C. Van der Pol oscillator (target in solid black) generated by the bias-learning RNN for a recurrent gain of 1 (dashed orange; see panel D) and a gain of 0.9 (dashed dark orange). Output represents the oscillator’s position, rescaled to [-1, 1]. D. (Left) Sensitivity to distribution of recurrent weights. The fully-trained and bias-learning networks had the same number of learnable parameters. Initial recurrent weight matrix had elements sampled from $(g/\sqrt{m})\mathcal{N}(0,1)$, where $g$ is the gain (Gain recurrent init.). Error bars denote SEM for $n=10$. (Right) Schematics of the fully-trained (top) and bias-learning (bottom) autonomous RNNs. Colored links denote trained weights.
@@
We next investigated whether bias learning relied on the statistics of the fixed recurrent weights. In light of Fig. 3B, we thus hypothesized that bias learning would be affected by changes in the weight distribution, because bias learning can only control the derivative. We initialized an i.i.d. Gaussian distributed weight matrix $W_{ij}\sim{\cal N}(0,g^{2}/N)$, where $g$ is referred to as its ‘gain’. We then trained bias-learning networks to generate a van der Pol oscillator (Fig. 3C). We found that bias learning required a large enough gain (at least $g=1$) and failed for $g<1$ (Fig. 3D). This was not purely due to a restricted dynamic range for the network activity since the network was able to reproduce the first peak of the oscillator and then flatlined (Fig. 3C). In contrast, fully-trained networks with the same number of training parameters (Fig. 3D) were not sensitive to the value of the gain at initialization. This result thus highlights that, when the hidden-layer size is fixed, the initial distribution of weights limits the capability of bias-learning networks.
3.4 Bias learning non-autonomous dynamical systems with RNNs
To further test bias learning in RNNs, we trained a RNN to predict future time-steps of a partially observed dynamical system, namely a single dimension of the Lorenz attractor (see Appendix §C.4for details). As in the autonomous DS case, only the biases of the input layer were trained and the weights were random and frozen. However, here the network received the observed dimension of the Lorenz system as an input. Given the observed dimension at time-point $t$, and its value at previous time-steps encoded in the RNN’s hidden state, the objective of the task was to predict the future value at $t+\tau$, where $\tau=27$ was chosen to be the half-width at the half-max of the auto-correlation of the observed dimension of the Lorenz system.
Figure 4: Learning non-autonomous dynamical systems. A. Validation $R^{2}$ vs. number of trainable model parameters for fully-trained (blue) and bias-learned (orange) RNNs. Training RNNs with bias learning became unstable below a network width of $64$. B. Predictions from the fully-trained and bias-learned networks (both with a hidden layer width of 1024) on a trajectory of the Lorenz system unseen during training. Standard deviation error bars were computed over 5 seeds, but are not visible due to their small magnitudes. C. Predictions of both the fully-trained and bias-learned networks diverge from the ground truth signal when one starts feeding back their own outputs as their inputs, in place of the ground-truth time-series (self-sustained, starting from the grey line).